hello my name is Frank
7-(2x-1)=10
is the next step
-14x-7=10
then add 7 to both sides
-14x=17

Answer by Algebraic(50)

Original equation: 7-(2x-1)=10
Step 1: You're going to want to begin by distributing the negative (-) symbol to everything in the parenthesis.
7 - (2x-1) = 10
7 - 2x + 1 = 10 (Remember: A negative and a positive sign make a negative sign (-), and a negative sign and a negative sign make a positive sign (+))
Step 2: Combine like terms
7 - 2x + 1 = 10
Add the 7 and 1 together.
8 - 2x = 10
Step 3: Subtract the positive 8 from both sides so you can get '-2x' by itself.
[(8 - 8) - 2x = (10 - 8)]
You'll get: -2x = 2
Step 4: Divide '-2' on both sides so you can get the 'x' by itself, like so:
-2x = 2
(-2/2) is 1, so 'x' stays by itself, and (2/-2) is a -1, so:
Step 5: The answer you'll get is x = -1